A stretched string is 1.97 m long and has a mass of 20.9 g . When the string is oscillated at 440 Hz , which is the frequency of the standard A pitch that orchestras tune to, transverse waves with a wavelength of 16.5 cm travel along the string. Calculate the tension in the string.
Answer #1

v= f*
v= 440x0.165
v= 72.6 m/s

v= (T/µ)0.5
72.6 = (T/(0.0209/1.97))0.5

T = 55.9 N.